# Unveiling galaxy chemical enrichment mechanisms out to cosmic dawn from direct determination of O & Ar abundances from JWST/NIRSPEC spectroscopy

### v1

@@ 2.3 Abundance determination
Emission lines fluxes for each galaxy are de-reddened using the calculated c(H), see Table 1, and their temperatures and densities are calculated using an iterative process from the relevant diagnostic lines using NEAT (see Wesson et al., 2012, section 3.3). For our observations, NEAT utilizes the temperature-sensitive [OIII] 4363 Å line, or the [OII] 7322,7332 Å lines for the two Keck/MOSFIRE observed galaxies, and the density-sensitive [OII] 3726,3729 Å and [SII] 6717,6731 Å doublets to obtain temperature and electron density for each galaxy spectra. O and Ar ionic abundances are measured from the observed fluxes of the O ([OII] 3726,3729 Å , [OIII] 4363,4959,5007 Å) and Ar ([ArIII] 7136 Å and/or [ArIV] 4711,4740 Å) lines respectively. ICF correction for O is negligible when lines pertaining to both O2+ (i.e, [OIII] 5007, 4959, 4363 Å) and O+ ([OII] 3727, 3729 Å) are observed. Elemental Ar abundances are obtained from the ionic abundances utilising the latest ionisation correction factors (ICF; Amayo et al. 2021) for Ar abundance measurement. The impact of the assumed ICF on the determined Ar abundances and thereby the implications for our results are discussed in Appendix C. Uncertainties are propagated through all steps of the analysis into the final values. The galaxy stellar masses, specific-star-formation rates (sSFR) and star-formation timescales (tSF) are discussed in Appendix D, and these along with the O and Ar abundances are presented in Table 2.
Name | z | log(M∗/M⊙) | log(sSFR) [yr-1] | tSF [Myr] | 12 + log(O/H) | 12 + log(Ar/H) | log(O/Ar) |
ERO 10612 | 7.66 | ||||||
CEERS 1536 | 5.038 | ||||||
GLASS 150029a | 4.584 | ||||||
CEERS 1665a | 4.488 | ||||||
CEERS 1651 | 4.382 | ||||||
GLASS 40066 | 4.02 | ||||||
JADES 19519 | 3.604 | ||||||
CEERS 11088 | 3.302 | ||||||
Q2343-D40b | 2.963 | – | – | – | |||
CEERS 3788 | 2.295 | ||||||
COSMOS 19985b | 2.188 | ||||||
COSMOS 20062b | 2.185 | ||||||
SGAS1723+34b | 1.329 |
Note. — Column 1: Name of galaxy; Column 2: galaxy redshift; Column 3: Measured Balmer decrement; Columns 4–8: Measured flux of different emission lines, relative to H=100, S/N is mentioned in brackets; Column 9: Source of publicly available flux and wavelength calibrated 1D JWST/NIRSPEC grating spectra; Column 10: Reference for first determination of O abundance and source for other line-fluxes of interest.
aIdentified from broad H lines as AGN host by Harikane et al. (2023).
bMikulski Archive for Space Telescopes (https://mast.stsci.edu).
cCEERS DR07 is available at https://ceers.github.io/dr07.html.
Note. — Column 1: Name of galaxy; Column 2: galaxy redshift; Column 3-5: Estimated mass, sSFR, and star-formation timescale of galaxies, from their SED fitted to broad-band photometry (see Appendix D); Columns 6–8: Estimated elemental abundances.
aIdentified from broad H lines as AGN host by Harikane et al. (2023).
bFor these galaxies, the JWST/NIRSPEC spectra were not available publicly. The abundances were estimated from line fluxes published in the literature.
2.4 Comparison with published abundance determinations
@@
In Figure 2 we show that our O abundances, determined following a uniform methodology, are consistent within error with previously published values. Our newly determined 12 + log(O/H) values have a mean offset of dex and standard deviation of dex compared to the literature values, well within the errors for individual determinations. We report in this work the first determination of Ar abundances and thereby log(O/Ar) for all these galaxies, except for SGAS1723+34 and Q2343-D40. For Q2343-D40, Rogers et al. (2024) reported an Ar abundance slightly lower than our determination, a potential consequence of their choice for a different ICF scheme (Izotov et al., 2006) than the one utilised in this work, or arising from differences in their abundance determination methodology. For SGAS1723+34, Welch et al. (2024) report an Ar abundance nearly identical to our determination (see Figure 2 [inset]).
@@
Figure 3a shows the position of the galaxies at z1.3–7.7 in the log(O/Ar) vs 12 + log(Ar/H) plane, and represents their state of chemical enrichment.
3.1 Sources at z1.3–3.4
@@
Six SFGs (z1.3–3.4; see Figure 3a) are consistent within error with the Milky Way (MW) solar neighbourhood Galactic Chemical Evolution (GCE) model (Kobayashi et al., 2020a), where CCSNe (including hypernovae) and SNe Ia dominate. Additionally, their position is consistent with the locus traced by the mean log(O/Ar) as a function of 12+log(Ar/H) for the higher mass (M M⊙) low-redshift (z) starbursts (log(sSFR)) from SDSS (Brinchmann et al., 2004, Bhattacharya et al. in prep). This remarkably shows that, just like the MW ISM (Matteucci, 2021; Kobayashi et al., 2020a) and higher-mass starbursts at z, the SFGs in our sample out to z3.4 are consistent with having chemical enrichment being driven mainly by CCSNe and SNe Ia, with the same nucleosynthesis yields and initial mass functions.
@@
The only other SFG (KBSS-LM1, z=2.396, Steidel et al. 2016) with direct auroral-line measurements and full-spectral UV-fitting222Other SFGs with [Fe/H] determined from full-spectral UV-fitting (e.g. Cullen et al., 2019, 2021; Stanton et al., 2024) have [O/H] (in lieu of abundance) determined from strong-line methods from optical emission-lines, which may be over-estimated if not calibrated against direct methods (Maiolino & Mannucci, 2019). has [O/Fe] at [Fe/H]-1.6. This is coincident with the MW solar neighbourhood GCE model in the [O/Fe] vs [Fe/H] plane (see Fig 3 in Kobayashi et al. 2020a), and in extension consistent with the z1.3–3.4 SFGs presented in this work. This shows, at least for some of our small sample of galaxies, that MW-like chemical enrichment sequences and their underlying mechanisms hold up to z1.3–3.4.
3.2 Sources at z3.6–7.7
Five SFGs (z3.6–7.7) have log(O/Ar) values below even the sequence traced by lower-mass (M M⊙) z starbursts (Figure 3a). The Ar abundance may be underestimated if a large fraction of Ar is condensed in dust grains but given that O and Ar have similar dust condensation temperatures (Savage & Sembach, 1996), it is unlikely that the lower log(O/Ar) for these SFGs is caused by more O being ejected with dust grains than Ar. This implies that the z3.6–7.7 SFGs are characterised by unusually high Ar production, relative to O. In the following subsections, we discuss possible scenarios that may be responsible for the low log(O/Ar) values for these galaxies.
@@
3.3 Potential chemical enrichment from intermittent starbursts at z3.6–7.7
@@
Keeping the same assumption of the chemical enrichment recipe for the solar neighbourhood (Kobayashi et al., 2020a), we construct a tailored GCE model considering two bursts of star-formation at very early times (with substantial infall of pristine gas between the bursts), to try and explain the position of the z3.6–7.7 SFGs in this plane (see Figure 3b). An intermittent star-formation model has previously been invoked to explain measurements of emission line fluxes in similar high redshift SFGs (Kobayashi & Ferrara, 2024). Note that the purpose of constructing this specific model is to check if the low log(O/Ar) values of these galaxies may be explained by considering an extreme star-formation history but keeping the MW-like CCSNe and SNe Ia dominated nucleosynthesis.
@@
In Figure 3b (inset), the star-formation history of the GCE model is plotted. An initial burst of star formation at the break of cosmic dawn is followed by a quiescent phase up to 1.1 Gyr after the birth of the universe, when the ISM is enriched through continued explosions of Type Ia SNe to very high 12 + log(Ar/H) values and minimum log(O/Ar) values (Figure 3b). This is followed by an infall of primordial gas that strongly dilutes the ISM, reducing 12 + log(Ar/H) keeping log(O/Ar) constant, and induces another episode of star-formation that starts re-enriching the ISM (Figure 3b inset). Thus a loop in the log(O/Ar) vs 12 + log(Ar/H) plane is formed which is a signature of gas infall (Spitoni et al., 2019; Arnaboldi et al., 2022; Kobayashi et al., 2023). Other models with less extreme star-formation histories (lower star formation rates with continuous or intermittent star-formation with and without gas infall) will occupy the parameter space within the loop spanned by this extreme model in this plane.
The model can successfully explain the positions of three of the z3.6–7.7 SFGs that have log(O/Ar). However, two of our SFGs that have log(O/Ar) values 3 lower than the MW solar neighbourhood GCE model (highlighted in Figure 3b) are also below the log(O/Ar) values reached by this model. These two galaxies are ERO 10612, which is the highest redshift galaxy in our sample at z=7.66, i.e, just 672 Myr after the birth of the universe, and CEERS 1651 at z=4.382 (1.37 Gyr after the birth of the universe). It is unclear whether CCSNe and SNe Ia dominated chemical enrichment models could explain these low-log(O/Ar) galaxies.
3.4 Additional potential sources of chemical enrichment at z3.6–7.7
@@
A potential mechanism for log(O/Ar)-poor stellar populations would be to assume a higher SNIa rate (e.g. with a higher binary fraction) relative to CCSNe for a given stellar population mass in the early universe, than is seen for the MW solar neighbourhood.
@@
Another potential rapid metal enrichment source is pair-instability supernovae (PISNe; Heger & Woosley, 2002; Nomoto et al., 2013) that have been predicted to evolve from massive ( M⊙) population III stars333First generation metal-free stars formed from pristine gas. having only Myr (Takahashi et al., 2018) lifespans. Their metal contribution to the ISM is expected to be visible only for a short time after the first generation of stars formed (Hartwig et al., 2018; Vanni et al., 2023), to be washed-out rapidly (Ji et al., 2015) once CCSNe enrichment processes begin after Myr (Kobayashi et al., 2020a). Given the small, albeit uncertain, tSF values (see Table 2) for ERO 10612 and CEERS 1651, PISNe may potentially be responsible for their observed Ar enhancement, but only if there has been no mixing with any pre-enriched ISM or mass-loss from pre-existing stars. From stellar nucleosynthesis yields of PISNe (Takahashi et al., 2018), the log(O/Ar) values of ERO 10612 and CEERS 1651 (see Figure 3b) are consistent with that expected from lower-mass (170-180 M⊙) and higher-mass (260-280 M⊙) PISN progenitors respectively444Computed assuming the observed O abundances of these two SFGs (see Table 2) and the log(O/Ar) values for non-rotating PISNe of different progenitor masses in Takahashi et al. (2018, see their Figure 9).. However, in GCE models with reasonable assumptions, the PISN enrichment causes the rapid decrease of log(O/Ar) to very low values at much lower metallicities (Kobayashi et al. in prep).
@@ 4 Summary and conclusions
We extend the use of the log(O/Ar) vs 12 + log(Ar/H) plane (Arnaboldi et al., 2022; Kobayashi et al., 2023) for inferring the mechanisms that govern galaxy chemical enrichment to SFGs, offering a direct analogy to the [/Fe] vs [Fe/H] plane for stars. We robustly obtain line-flux measurements for 7 SFGs from their flux and wavelength calibrated 1D JWST/NIRSPEC spectra (see Table 1), in addition to 4 SFGs where such flux measurements were available from their literature sources (see Section 2). We then directly determine O and Ar abundances for these 11 SFGs at z1.3–7.7 from observations of temperature sensitive auroral lines (Table 2).
We present their positions in the log(O/Ar) vs 12 + log(Ar/H) plane (Figure 3a). Six SFGs (z1.3–3.4) are consistent within error with a MW-like CCSNe and SNe Ia dominated chemical enrichment model (Kobayashi et al., 2020a). Thus at least for some of our small sample of galaxies, MW-like chemical enrichment sequences and their underlying mechanisms are in place as early as z1.3–3.4.
Five SFGs (z3.6–7.7), however, are log(O/Ar)-poor compared to the MW GCE model track (Figure 3a). The log(O/Ar) values of three of these may be explained through a tailored GCE model with early intermittent star-formation, but keeping the MW-like CCSNe and SNe Ia dominated chemical enrichment (see Section 3.3). For the other two, ERO 10612 and CEERS 1651, additional potential sources of chemical enrichment may be required. Tailored GCE models will be utilised to better understand the rapid enrichment of these SFGs (Kobayashi et al. in prep).
@@
Appendix A Flux calibration and line detection from JWST/NIRSPEC spectra
As the wavelength and flux calibrated spectra were downloaded from public archives, one cannot rule out possible significant differences of the line-fluxes and abundances determined from them in this work with those previously published (references noted in Table 1). In Figure 2 we show that our determined O abundances are always consistent within error with previously published values, showing negligible offset ( dex) and dispersion ( dex, see Section 2.4). Our measured [OIII] 4363 Å line fluxes (see Table 1) are also consistent with their previously published values. The tight correlation between the O abundance values from literature and those determined in this work implies that the flux calibration in the spectra range 3700–6750 Å is robust and consistent between the spectra used here and those utilised previously. Since the [ArIII] 7136 Å line is sufficiently close to the redder limit of the aforementioned wavelength range, the flux calibration properties apply also to this line.
We carried out additional tests to verify the reliable detections of the [OIII] 4363 Å and [ArIII] 7136 Å line line fluxes in the publicly available JWST/NIRSPEC spectra. If the observed rest-frame 1D spectra (normalised to a given line flux) are stacked, the S/N of a detected line should show an increase of an order of the square root of the number of spectra being stacked, in case of poissonian noise (e.g. Arnaboldi et al., 2002). We make two stacked SFG spectra with the flux and wavelength calibrated 1D JWST spectra (each normalised to H flux), one for z with CEERS 11088 and 3788 (as we do not have the publicly available spectra for other SFGs at these redshifts; see Table 2), and another for z with CEERS 1536 and 1651, GLASS 40066 and JADES 19519 (as only these four are SFGs that have the [ArIII] 7136 Å line). For the z stack, we find that the [OIII] 4363 Å and [ArIII] 7136 Å lines are detected with S/N of 5.97 and 7.3 respectively, showing the expected increase in S/N compared to the line detection in individual spectra. For the z stack, we find that the [OIII] 4363 Å and [ArIII] 7136 Å lines are detected with S/N of and respectively, also showing the expected S/N increase compared to the detection in individual spectra.
We additionally note that the CEERS 11088 galaxy was observed with the [ArIII] 7136 Å line detected in two separate JWST/NIRSPEC grism observations (G235M/F170LP and G395M/F290LP) as it occurs in the overlapping wavelength region. The [ArIII] line is detected with S/N in both spectra from different grism observations, with similar line flux values within error, as reported in Table 1. The independent line detection and concordance of measured fluxes additionally certifies the [ArIII] 7136 Å line detection for this galaxy but also attests to the reliable flux calibration for those galaxies with CEERS DR07 spectra (see Table 1). Note also that while we use the flux and wavelength calibrated 1D spectra from the source reported in Table 1, we have also checked other sources when available, such as the DAWN JWST archive (https://dawn-cph.github.io/dja/) that employ different flux and wavelength calibrations. While there may be minor discrepancy in the measured fluxes of individual lines when comparing the calibrated 1D spectra of a galaxy from different public sources, we find that, for the galaxies in our sample, the determined O and Ar abundances are always consistent within error.
Appendix B Impact of line-blending on Ar abundances

### v2

@@ 2.4 Abundance determination
Emission lines fluxes for each galaxy are de-reddened using the calculated c(H), see Table 2, and their temperatures and densities are calculated using an iterative process from the relevant diagnostic lines using NEAT (see Wesson et al., 2012, section 3.3). For our observations, NEAT utilizes the temperature-sensitive [OIII] 4363 Å line, or the [OII] 7322,7332 Å lines for the two Keck/MOSFIRE observed galaxies, and the density-sensitive [OII] 3726,3729 Å , [ArIV] 4711,4740 Å and [SII] 6717,6731 Å doublets to obtain temperature and electron density for each galaxy spectrum. For three galaxies (see Table 2), we do not observe the required doublets to determine electron densities. In these cases, 1000 cm-3 is adopted, however the value is expected to have negligible impact on the determined abundances as the emissivities of the auroral line transitions are nearly density independent at the low densities determined for our observed sources (e.g. Ferland et al., 2013). Assuming a lower electron density of 200 cm-3, similar to many galaxies in our sample (see Table 2) leads to consistently lower log(O/Ar) values of 0.05 dex and higher 12 + log(Ar/H) values by 0.03 dex, well within the estimated errors.
@@
O and Ar ionic abundances are measured from the observed fluxes of the O ([OII] 3726,3729 Å , [OIII] 4363,4959,5007 Å) and Ar ([ArIII] 7136 Å and/or [ArIV] 4711,4740 Å) lines respectively. Ionisation correction factor (ICF) for O is negligible when lines pertaining to both O2+ (i.e, [OIII] 5007, 4959, 4363 Å) and O+ ([OII] 3727, 3729 Å) are observed. Elemental Ar abundances are obtained from the Ar2+ ionic abundances utilising the ICF from Amayo et al. (2021) when [ArIII] 7136 Å is detected. Only for ERO 10612, the reported Ar abundance is the Ar3+ ionic abundance obtained from [ArIV] 4711,4740 Å observations with no ICF correction. The impact of the assumed ICF on the determined Ar abundances and thereby the implications for our results are discussed in Appendix C. Uncertainties are propagated through all steps of the analysis into the final abundance values. Along with the O and Ar abundances in Table 2, we report the galaxy stellar masses, specific-star-formation rates (sSFR) and star-formation timescales (tSF), which are also are discussed in Appendix D.
2.5 Comparison with previously published abundance determinations
In Figure 3 we show our determined O abundances against previously published values. Most of our galaxies (Curti et al. 2023: ERO 10612; Sanders et al. 2024: CEERS 1665, Isobe et al. 2023: GLASS 40066; Morishita et al. 2024: JADES 19519; Rogers et al. 2024: Q2343-D40; Sanders et al. 2023: COSMOS 19985, COSMOS 20062; Welch et al. 2024: SGAS1723+34) are consistent within the errors, but some offset is observed for a few others (Nakajima et al. 2023: ERO 10612, CEERS 1536, GLASS 150029; Sanders et al. 2024: CEERS 1651, CEERS 11088, CEERS 3788). Our newly determined 12 + log(O/H) values have a mean offset of dex and standard deviation of dex compared to the literature values. The small differences are mainly related to differences in flux and wavelength calibration between our utilised archival spectra and those utilised by previous studies. This is illustrated for ERO 10612 where our determined O abundance is consistent with that reported by Curti et al. (2023) but not with that reported by Nakajima et al. (2023), with both these authors using different in-house spectral flux and wavelength calibrations. Indeed, applying our flux measurement and abundance determination procedure to older versions of JWST NIRSPEC spectra for our studied galaxies, as released by previous authors333Particularly public versions of the spectra for CEERS DR0.7 and JADES DR3 were analysed., we find our determined 12 + log(O/H) values have a lower mean offset of dex and lower standard deviation of dex compared to literature works.
@@ 2.5 Comparison with previously published abundance determinations
We report in this work the first determination of Ar abundances and thereby log(O/Ar) for all these galaxies except SGAS1723+34 and Q2343-D40. For Q2343-D40, Rogers et al. (2024) reported an Ar abundance slightly lower than our determination, a potential consequence of their choice for a different ICF scheme (Izotov et al., 2006) than that utilised here, or arising from differences in their abundance determination methodology. For SGAS1723+34, Welch et al. (2024) report an Ar abundance nearly identical to our determination (see Figure 3 [inset]).
@@
Figure 4 [Left] shows the position of the galaxies at z1.3–7.7 in the log(O/Ar) vs 12 + log(Ar/H) plane, and represents their state of chemical enrichment.
@@
Seven SFGs (z1.3–4; see Figure 4 [Left]) are consistent within error with the Milky Way (MW) solar neighbourhood Galactic Chemical Evolution (GCE) model (Kobayashi et al., 2020a), where CCSNe (including hypernovae) and SNe Ia dominate in a self-regulated scenario with no inflows or outflows. Additionally, their position is consistent with the locus traced by the mean log(O/Ar) as a function of 12+log(Ar/H) for the higher mass (log(M∗/M⊙) = 9.41) low-redshift (z) starbursts from SDSS (Brinchmann et al., 2004, Bhattacharya et al. in prep). This remarkably shows that, just like the MW ISM (Matteucci, 2021; Kobayashi et al., 2020a) and higher-mass starbursts at z, the SFGs in our sample out to z4 are consistent with having chemical enrichment being driven mainly by CCSNe and SNe Ia, with the same nucleosynthesis yields and initial mass functions.
We note that O & Ar abundance determination from JWST/NIRSPEC spectra have now been reported independently for an additional 8 galaxies at z1.8–5.2 by Stanton et al. (2024) and 4 more at z3.2–4.7 by Stiavelli et al. (2024). Almost all these galaxies have log(O/Ar) and 12+log(Ar/H) values consistent with the MW GCE model (see also Figure 5 in Stanton et al. 2024).
@@
The only other SFG (KBSS-LM1, z=2.396, Steidel et al. 2016) with direct auroral-line measurements and full-spectral UV-fitting444Other SFGs with [Fe/H] determined from full-spectral UV-fitting (e.g. Cullen et al., 2019, 2021; Stanton et al., 2024) have [O/H] (in lieu of abundance) determined from strong-line methods from optical emission-lines, which may be over-estimated if not calibrated against direct methods (Maiolino & Mannucci, 2019). has [O/Fe] at [Fe/H]-1.6. This is coincident with the MW solar neighbourhood GCE model in the [O/Fe] vs [Fe/H] plane (see Fig 3 in Kobayashi et al. 2020a), and in extension consistent with the z1.3–4 SFGs presented in this work.
On the other hand, four SFGs (z3.5–7.7) have log(O/Ar) values below the MW GCE model considering also their errors and scatter, and even below the sequence traced by lower-mass (M M⊙) z starbursts (Figure 4 [Left]). O and Ar have similar dust condensation temperatures (Savage & Sembach, 1996), therefore it is unlikely that their log(O/Ar) value is underestimated on account of preferential ejection of only condensed O with dust grains.
Therefore in our small sample of galaxies MW-like self-regulated chemical enrichment sequences and their underlying mechanisms hold up to z4. At z3.5 or so, SFGs appear to be characterised by higher Ar abundance relative to O. In the following subsections, we discuss possible scenarios that may be responsible for the low log(O/Ar) values for the higher redshift galaxies in our sample.
@@
3.1 Potential chemical enrichment from intermittent starbursts at high redshift
@@
Keeping the same assumption of the chemical enrichment recipe for the solar neighbourhood (Kobayashi et al., 2020a), we construct an illustrative GCE model considering two bursts of star-formation at very early times (with substantial infall of pristine gas between the bursts), to try and explain the position of the four low log(O/Ar) SFGs in this plane (see Figure 4 [Right]). An intermittent star-formation model has previously been invoked to explain measurements of emission line fluxes in similar high redshift SFGs (Kobayashi & Ferrara, 2024). Note that the purpose of constructing this specific model is to investigate whether the low log(O/Ar) values of these galaxies may be explained by considering an extreme star-formation history while keeping the MW-like CCSNe and SNe Ia dominated nucleosynthesis.
@@
In Figure 4 [Right] (inset), the star-formation history of the GCE model is plotted. An initial burst of star formation at the break of cosmic dawn is followed by a quiescent phase up to 1.1 Gyr after the birth of the universe, when the ISM is enriched through continued explosions of Type Ia SNe to very high 12 + log(Ar/H) values and minimum log(O/Ar) values. This is followed by an infall of primordial gas that strongly dilutes the ISM, reducing 12 + log(Ar/H) keeping log(O/Ar) constant, and induces another episode of star-formation that starts re-enriching the ISM. Thus a loop in the log(O/Ar) vs 12 + log(Ar/H) plane (Figure 4 [Right]) is formed which is a signature of gas infall (Spitoni et al., 2019; Arnaboldi et al., 2022; Kobayashi et al., 2023). Other models with less extreme star-formation histories (lower star formation rates with continuous or intermittent star-formation with and without gas infall) will occupy the parameter space within the loop spanned by this extreme model in this plane. The model can successfully explain the positions of three of the SFGs at z3.5–5 that have log(O/Ar). Note that SFGs are not expected to appear on the constant log(O/Ar)2.1 line, except for at very low 12+log(Ar/H) values, for the model (Figure 4 [Right]) as this signifies the rapid dilution of the ISM and star-formation only follows afterwards.
We note however that ERO 10612, which is the highest redshift galaxy in our sample at z=7.66, i.e, just 672 Myr after the birth of the universe555We assume Planck cosmological parameters (Planck Collaboration et al., 2020): Hubble constant H0 = km s-1 Mpc-1, matter density parameter = . With these assumptions we determine the age of the universe at given redshifts for the galaxies in our sample in Figure 4 [Right]. We assume epoch of first star formation at z=25, within the expected z=20–30 range (Bromm et al., 2009). This has been applied to the GCE model shown in Figure 4 [Right] (inset)., has a log(O/Ar) value 1.8, that is below the log(O/Ar) values reached by the intermittent star-formation model (see Figure 4). Additional physical mechanisms, such as outflows, may be at play for this galaxy. A possible scenario may be one where after the first burst of star-formation the CCSNe ejecta with O and Ar are expelled in an outflow while Ar produced over longer timescales is retained, eventually reaching the observed log(O/Ar) value 1.8.
It remains to be demonstrated whether CCSNe and SNe Ia dominated chemical enrichment models could explain the low-log(O/Ar) value for ERO 10612, especially given the short timespan after the birth of the universe within which such abundance values need to be reached. Future GCE simulations for this galaxy exploring different star-formation histories with inflows and outflows may address this property. Additional sources of chemical enrichment described in literature sources may also potentially be at play, as discussed concisely in the following section.
3.2 Additional potential sources of chemical enrichment at high redshift
@@
A potential mechanism for additional Ar production leading to log(O/Ar)-poor stellar populations would be to assume a higher SN Ia rate (e.g. with a higher binary fraction) relative to CCSNe for a given stellar population mass in the early universe, than is seen for the MW solar neighbourhood.
@@
Another potential rapid metal enrichment source is pair-instability supernovae (PISNe; Heger & Woosley, 2002; Nomoto et al., 2013) that have been predicted to evolve from massive ( M⊙) population III stars (first generation metal-free stars formed from pristine gas) having only Myr (Takahashi et al., 2018) lifespans. However, their metal contribution to the ISM is expected to be visible only for a short time after the first generation of stars formed (Hartwig et al., 2018; Vanni et al., 2023), to be washed-out rapidly (Ji et al., 2015) once CCSNe enrichment processes begin after Myr (Kobayashi et al., 2020a). Given the small, albeit uncertain, tSF value of ERO 10612 (see Table 2), PISNe could in principle explain its low log(O/Ar) values based on model PISN O and Ar yields (Takahashi et al., 2018), but only if there has been no mixing with any pre-enriched ISM or mass-loss from pre-existing stars. This possibility underlines the interest in further detailed studies of ERO 10612 and other potentially similar systems. However, in GCE models with conventional assumptions, the PISN enrichment causes the rapid decrease of log(O/Ar) to very low values at much lower metallicities (Kobayashi et al. in prep).
@@ 4 Summary and conclusions
We extend the use of the log(O/Ar) vs 12 + log(Ar/H) plane (Arnaboldi et al., 2022; Kobayashi et al., 2023) for inferring the mechanisms that govern galaxy chemical enrichment to SFGs, offering a direct analogy to the [/Fe] vs [Fe/H] plane for stars. We robustly obtain line-flux measurements for eight SFGs from their flux and wavelength calibrated 1D JWST/NIRSPEC spectra (see Table 1), in addition to three SFGs where such flux measurements were available from their literature sources (see Section 2). We then directly determine O and Ar abundances for these 11 SFGs at z1.3–7.7 from observations of temperature sensitive auroral lines (Table 2).
We present their positions in the log(O/Ar) vs 12 + log(Ar/H) plane (Figure 4a). Seven SFGs (z1.3–4) are consistent within error with a MW-like CCSNe and SNe Ia dominated chemical enrichment model (Kobayashi et al., 2020a). Four SFGs (z3.5–7.7) are found to be log(O/Ar)-poor compared to the other seven aforementioned SFGs and the MW GCE model track (Figure 4).
Thus, in the majority of our small sample of galaxies, MW-like self-regulated chemical enrichment sequences and their underlying mechanisms may be in place as early as z1.3–4. This is corroborated for independent galaxy samples in Stanton et al. (2024) and Stiavelli et al. (2024).
The low log(O/Ar) values of three SFGs at higher redshift (z3.5–5.) may be explained through a tailored GCE model with early intermittent star-formation, but keeping the MW-like CCSNe and SNe Ia dominated chemical enrichment (see Section 3.1). For ERO 10612 (z=7.66), deeper observations would improve the S/N of the faint [ArIV] 4740 Å line. Exploration of different star-formation histories, potentially with more bursts, considering also inflows and outflows that preferentially eject O and/or additional potential sources of Ar enrichment may be required to explain the low log(O/Ar) determined for these galaxies. Dedicated GCE models will be utilized to better understand the rapid Ar enrichment of these SFGs (Kobayashi et al. in prep).
@@
Appendix A Reliability of flux and wavelength calibrations for JWST/NIRSPEC spectra
The spectra are flux calibrated with an approximately 15% absolute flux accuracy, while field-dependent variations may be as large as 10% as per the latest JWST/NIRSPEC MOS calibration pipeline. Emission line fluxes estimated from medium resolution grating spectra seem to consistently have 10% higher flux values compared to PRISM spectra and NIRCAM photometry (Bunker et al., 2023), although line flux ratios (as used here) seem unaffected. As per the same latest calibrations, wavelength calibration is accurate till 15 km/s and 40 km/s for high and medium resolution gratings respectively. Detector artifacts have been more effectively removed by the pipeline for the latest available spectra.
Appendix B Impact of line-blending on Argon abundances